PMC lost lock again at around 16:00 April 12. I was able to lock it again but the transmission is only 0.6 now and REFL is 0.14.
Rana came in and realigned the PMC stirring mirrors. Now the transmission is 0.757V, and the REFL is 0.03V.
I noticed that the PZT was around 250V. Given that the PMC got unlocked at 16:00, which is around the peak temperature time in the lab (lagging behind the outside weather), due to the PZT voltage going down to 0V, I figured that the PZT voltage would go up during the night when the lab gets cold and therefore will likely go out of range again.
I found a different working point at 150V and relocked the PMC.